OpenStudy (noahbred):

Help for medals 27 divided by 1/3

OpenStudy (smilie2015):

when does 27 comes in 3 table ?

OpenStudy (smilie2015):

3*9=?

OpenStudy (noahbred):

27

OpenStudy (smilie2015):

soo 27/3 = ?

OpenStudy (smilie2015):

ohhh sorry 27 is divided by 1/3 \[27\div 1/3\]

OpenStudy (md152727):

I can help

OpenStudy (smilie2015):

27* 3= ?

OpenStudy (md152727):

81

OpenStudy (noahbred):

81

OpenStudy (md152727):

so 1/3=0.3333333333333

OpenStudy (smilie2015):

yup :)

OpenStudy (md152727):

any multiple choices?

OpenStudy (noahbred):

yah

OpenStudy (md152727):

the answer ranges from 8.90-9

OpenStudy (md152727):

Can you please post them?

OpenStudy (noahbred):

A. 9 B. 81 C. –9 D. –81

OpenStudy (noahbred):

So Beeeee

OpenStudy (md152727):

What do you think?

OpenStudy (md152727):

ok, it is not C or D

OpenStudy (md152727):

umm

OpenStudy (md152727):

Yeah, B I am sure

OpenStudy (noahbred):

ty guuys